# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v1

@@ Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting the universal principles underlying languages. Particularly, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended to child languages and animal signals. However, the argument supporting this interpretation has not been empirically tested. To address this problem, this study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the implicit assumptions in the argument. Using English corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in parse trees, as well as in the PCFG that approximates these trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument to child languages and animal signals, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
11footnotetext: This work was conducted while KN and KK were at the University of Tokyo.
@@ 1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages, providing insights into the mechanisms that give rise to the properties. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 2016) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal principles of languages.
@@
This study focuses on the power-law decay of correlation. In languages, distant elements can remain strongly correlated. For example, subject–verb agreement in English can occur at arbitrarily large distances, and a single sentence can have semantic influence over an entire text. Previous studies have measured correlations between words, characters, phones, and other elements using metrics such as correlation functions and mutual information (MI). These studies have consistently found that the correlation decays according to a power-law function of the distance between elements (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024). This phenomenon indicates that the correlation is significant even at large distances.
@@
Moreover, the power-law decay of correlation has been interpreted as evidence supporting the linguistic view that languages are organized hierarchically in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This perspective has also been extended beyond adult human languages. Previous research has observed similar power-law behaviors in child languages and in animal signals, arguing that these may also possess underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
@@
This interpretation is justified by the argument of Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in structures and assumed that the sequential distance grows exponentially with the structural distance. These give rise to the power-law decay of correlation in sequences.
@@
However, this argument rests on several implicit assumptions. Previous studies have not tested their validity. To address this, we examine whether these assumptions hold for hierarchical structures in syntax, i.e., parse trees. Although parsing schemes vary in detail, syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables us to perform statistical analyses on treebanks to test these assumptions.
@@
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) Are PCFGs adequate models for the hierarchical structures of natural languages? To answer these questions, we analyze treebanks.
@@
As a result, our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures. Therefore, it is needed to reconsider the relationship between the power law and hierarchical structures.
@@
In addition, we perform the same analyses on trees generated by the PCFG that approximates syntactic structures, to compare them with natural languages. These analyses highlight the differences between PCFGs and natural languages. Furthermore, their statistical properties are also different from those assumed in the argument by Lin and Tegmark (2017). This means that the PCFG does not converge to its asymptotic behavior within the scale of a single sentence. Consequently, it is difficult to apply the proposed argument not only to single sentences, but also to other types of short sequences such as child languages and non-human signals.
To investigate if the argument holds beyond the sentence level, future research will need to formalize the description of hierarchical structures of entire texts, such as discourse. Otherwise, we need to seek for an alternative argument that account for the power-law behavior.
@@ 2.1 Power-law Decay of Correlation
In contrast, the power-law decay, $C\sim r^{-\alpha}$, is slower than the exponential one for any correlation length $\xi$, meaning that the correlation is significant for any long distances. In other words, the scale of correlation diverges. Moreover, power-law functions are scale-invariant. For the scale transformation $r^{\prime}=kr$, we can rescale the correlation by $C^{\prime}=k^{-\alpha}C$ so that the relation remains the same, $C^{\prime}\sim r^{\prime-\alpha}$. This property also implies the divergence of the scale.
Therefore, the power-law decay observed in various corpora means that the scale diverges in natural languages. Intuitively, a single element change can significantly influence the overall structure of a text.
@@ 2.2 Power Law and Hierarchical Structures
In the study of human languages and animal signals, the power-law decay of correlation has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs. PCFGs generate tree structures from the root to leaves iteratively, where the state of a parent node determines the states of its children. The leaves correspond to a sequence, while the whole tree is the underlying hierarchical structure. PCFGs are a probabilistic extension of context-free grammars (CFGs; Chomsky, 1956). CFGs were originally introduced to describe the nested structures, such as the if…then construction. Due to their simplicity and ability to capture fundamental aspects of hierarchical syntactic structures, PCFGs have been used to model syntactic structures (Charniak, 1997). Beyond linguistics, PCFGs have also been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
@@
We give an overview of the discussion by Lin and Tegmark (2017). To understand their discussion, it is useful to distinguish between sequential and structural distances, which we refer to as $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. The former is defined along the sequence. On the other hand, the latter is the distance in the hierarchical structure behind the sequence, as shown in Figure 1.
Lin and Tegmark (2017) have proven that the correlation asymptotically decays according to an exponential function of the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$. They also assumed that generated trees are balanced, i.e., they have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is asymptotically exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Therefore, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
@@
It is widely accepted among linguists that the syntactic structure of a sentence is hierarchical. Moreover, formal descriptions of these structures have been well established. Therefore, this study aims to examine the validity of the proposed argument, focusing on syntactic structures.
Meanwhile, the proposed argument is not limited to syntactic structures in principle. It may also apply to other types of structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchical organization may explain why the decay of correlation follows a power law at scales beyond the sentence level, extending across entire texts. The structures may also be more abstract. Previous research has observed the power-law decay in sequences without syntax, such as child languages and animal signals, arguing that these also have underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, the existence and the formalization of these types of hierarchical structures remain unclear. Therefore, we are currently unable to assess whether the proposed argument accounts for the power law at scales beyond the sentence level or in early-stage or non-human languages.
@@ 2.3 Rethinking Previous Research
The argument by Lin and Tegmark (2017) raises several important questions that must be addressed to determine whether it can apply to syntactic structures. First, do hierarchical structures in natural languages actually exhibit the behaviors assumed in their argument? More specifically:
@@
    Does the correlation in hierarchical structures decay exponentially with the structural distance?
@@
    Are the hierarchical structures sufficiently balanced such that the sequential distance grows exponentially with the structural distance?
Particularly, the latter appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth is slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
Figure 2: Example of a parse tree from the BLLIP corpus, preprocessed as described in Section 3.1. This tree has a strong right-branching bias.
Moreover, it is unlikely that PCFGs model syntactic structures sufficiently well. Linguistically, some languages cannot be described by CFGs (Shieber, 1985; Culy, 1985; Senuma and Aizawa, 2024). In addition, empirically, naive parsers based on PCFGs are insufficient for parsing (Johnson et al., 2006; O’Donnell et al., 2009). If syntactic structures deviate from PCFGs significantly, the discussion by Lin and Tegmark (2017) based on PCFGs should not be applied to them. This leads to the last question:
@@
    Do PCFGs describe the statistical properties of hierarchical structures sufficiently?
In this study, we perform statistical analyses on parse trees to address these questions.
@@ 3.1 Datasets and Preprocessing
We adopt the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000), a large-scale English treebank that provides Penn Treebank (Marcus et al., 1993)-style syntactic annotations for three years’ worth of Wall Street Journal articles. This treebank contains approximately 30 million words and one million sentences in total. In advance, we remove the phonological null elements such as traces of movement and ellipsis to ensure the alignment with the WikiText-103 corpus (Merity et al., 2017) analyzed in Appendix D, which does not include traces.
@@
The number of child nodes differs across nodes in the dateset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees using the Tree.chomsky_normal_form method from the NLTK library (Bird and Loper, 2004). We then apply the Tree.collapse_unary method from the same library to remove the unary branches produced by the former method. In addition, because a larger number of possible node states requires more data for reliable statistical estimation, we replace phrasal and part-of-speech (POS) tags with a smaller set of categories. Specifically, we refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. We present the eleven categories used in our analyses in Appendix A. An example of a preprocessed parse tree is shown in Figure 2.

### v2

@@ Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting universal mechanisms underlying languages. Particularly, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended to child speeches and animal signals. However, the argument supporting this interpretation has not been empirically tested in natural languages. To address this problem, the present study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the assumptions in the argument. Using English and Japanese corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in natural language parse trees, as well as in the PCFG that approximates these parse trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument to child speeches and animal signals, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
11footnotetext: This work was conducted while KN and $K$ were at the University of Tokyo.
@@ 1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages, providing insights into the mechanisms that give rise to the properties. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 1949) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal mechanisms underlying these power laws.
@@
This study focuses on the power-law decay of correlation. Previous studies have measured correlations between phones, characters, words, and other elements using metrics such as correlation functions and mutual information (MI), across various corpora. These studies have consistently found that the correlation decays according to a power-law function of the distance between elements (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024).
The power-law decay is qualitatively slower than the exponential decay, which is typically observed when interactions between elements are local, as in Markov chains. Therefore, the power-law decay of correlation in natural languages indicates that distant elements remain strongly correlated. This observation is consistent with linguistic intuition. For example, syntactic dependencies can, in principle, span arbitrarily long distances through repeated center-embedding, and a word introduced early in a context may still influence interpretation much later.
@@
Moreover, the power-law decay of correlation has been interpreted as evidence for the linguistic view that languages are organized hierarchically in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This interpretation has also been extended beyond speeches and written texts by human adults. Previous research has observed similar power-law behaviors in child speeches and in animal signals, arguing that these may also possess underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
@@
This interpretation is justified by Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in structures. In addition, they assumed that the sequential distance grows exponentially with the structural distance. These two give rise to the power-law decay of correlation in sequences.
However, several assumptions underlying this argument seem implausible in light of well-established linguistic facts. Yet, previous studies have not tested the validity of these assumptions quantitatively using empirical data.
@@
To address this, we examine whether these assumptions hold for hierarchical structures in natural language syntax, i.e., parse trees. Syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables us to perform statistical analyses on treebanks to test these assumptions.
@@
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in syntactic structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) To what extent do the statistical properties of syntactic structures deviate from those of PCFGs? To answer these questions, we analyze English and Japanese treebanks.
@@
As a result, our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures in natural languages. Therefore, it is necessary to reconsider the relationship between the two established facts, the power-law decay of correlation and the hierarchical syntactic structures in natural languages.
@@
In addition, we perform the same analyses on trees generated by the PCFG that approximates parse trees, to highlight differences between natural language syntactic structures and those generated by the PCFG. These analyses further show that the MI in structures generated by the PCFG does not exhibit exponential decay when the length of generated sequences is comparable to that of single sentences. Although Lin and Tegmark (2017) proved that the decay converges to an exponential form at sufficiently large distances, which their argument relies on, the decay follows a non-exponential form at distances typically observed in natural language parse trees. The parse trees are too small for this asymptotic exponential behavior to emerge. Consequently, it is difficult to apply their argument not only to single sentences by adults but also child speeches and non-human signals, in which the sequential distance is typically as short as in single sentences.
While the present study focuses on the hierarchical syntactic structures of individual sentences, the structures of entire texts consisting of multiple sentences, such as discourse structures, also appear to be hierarchical. The argument by Lin and Tegmark (2017), if applied to these structures, might account for the power-law decay of correlation in natural languages.
It would therefore be fruitful to investigate the relationship between the power-law behavior and hierarchical structures at the discourse level. Although previous studies have proposed formal frameworks for discourse annotation (Miltsakaki et al., 2004; Zeldes et al., 2025), the size of existing corpora remain too small to examine the validity of the argument. Developing sufficiently large-scale discourse corpora would be an important next step for future research.
Another direction is to explore alternative mechanisms that could give rise to power-law behavior in natural languages.
@@ 2.1 Power-law Decay of Correlation
Previous research has shown that the correlation in natural languages decays according to a power-law function of distance. This phenomenon has been observed across various corpora and languages (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024).
@@
In contrast, the power-law decay is expressed as $C\sim r^{-\alpha}$. Although the right-hand side resembles a polynomial form, the exponent $\alpha$ is a real number rather than a natural number. This decay is slower than the exponential one for any correlation length $\xi$, meaning that the correlation is significant for any long distances. In other words, the scale of correlation diverges.
Moreover, power-law functions are scale-invariant. For the scale transformation $r^{\prime}=kr$, we can rescale the correlation by $C^{\prime}=k^{-\alpha}C$ so that the relation remains the same, $C^{\prime}\sim r^{\prime-\alpha}$. This property also implies the divergence of the scale.
Therefore, the power-law decay of correlation observed in various corpora means that the scale diverges in natural languages. Intuitively, this implies that a local change can propagate and affect the global structure of a text.
@@ 2.2 Power Law and Hierarchical Structures
In studies of human languages and animal signals, this power-law behavior has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs.
PCFGs generate tree structures in a top-down manner, from the root to the leaves iteratively, where the category of a parent node determines the categories of its children. The leaves correspond to a surface sequence, while the whole tree represents the underlying hierarchical structure.
PCFGs are a probabilistic extension of context-free grammars (Chomsky, 1956), which were originally introduced to describe the nested structures, such as the if…then construction. PCFGs have also been used to model syntactic structures (Charniak, 1997), due to their simplicity and ability to capture fundamental aspects of hierarchical syntactic structures.
We provide an overview of the discussion in Lin and Tegmark (2017). To understand their discussion, it is useful to distinguish between sequential and structural distances, which we refer to as $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. The former is defined along the sequence. The latter is the distance in the hierarchical structure behind the sequence, as shown in Figure 1.
@@
Lin and Tegmark (2017) have proven that in PCFGs, the correlation decays exponentially with the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$, at sufficiently large distances. Note that the decay may follow non-exponential forms at shorter distances. They also assumed that generated trees are balanced, i.e., the trees have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is asymptotically exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Therefore, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
@@
The present study aims to examine the validity of this proposed argument. Throughout this study, we focus on syntactic structures. It is widely accepted in linguistics that these structures are hierarchical, with well-established formal descriptions. Using large-scale datasets based on this formalism, we can test if the proposed argument holds for syntactic structures.
It is worth noting that, in principle, PCFGs are models not only for syntactic structures but also for more general hierarchical structures. Indeed, PCFGs have been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
From this perspective, the argument by Lin and Tegmark (2017) may also apply to other types of structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchical organization may explain why the decay of correlation follows a power law beyond individual sentences, extending across entire texts.
However, it is difficult to empirically test whether the argument holds for multi-sentence texts. Although previous studies have proposed formal frameworks for annotating discourse structures (Miltsakaki et al., 2004; Zeldes et al., 2025), existing discourse corpora are not yet large enough to reliably estimate the statistical quantities introduced in Section 3, which we use to examine the validity of the assumptions in Lin and Tegmark (2017). Reliable estimation of these quantities requires substantial amounts of data.
@@ 2.3 Rethinking Previous Research
To test whether the argument by Lin and Tegmark (2017) holds for syntactic structures in natural languages, several important questions must be addressed.
First, do syntactic structures in natural languages actually exhibit the behaviors assumed in their argument? More specifically:
@@
    Does the correlation in syntactic structures decay exponentially with the structural distance?
@@
    Are syntactic structures sufficiently balanced such that the sequential distance grows exponentially with the structural distance?
Figure 2: Example of a parse tree from BLLIP, preprocessed as described in Section 3.1. This tree has a strong right-branching bias.
Particularly, the second one appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth can be slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
In addition, Lin and Tegmark (2017) demonstrate their argument using PCFGs as models of hierarchical structures in natural languages. However, the distribution that PCFGs can represent is highly limited. In this model, the distribution of a subtree depends only on its root, a property we refer to as the context-free independence. Previous studies have suggested that relaxing this property is necessary for modeling syntactic structures (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). Therefore, the last question is:
@@
    Do the statistical properties of syntactic structures deviate significantly from those of PCFGs?
These three questions remain unresolved. First, although extensive research has measured the correlation in texts and speeches, the correlation in the underlying structures have not been quantified. Therefore, the answer to question (i) remains elusive. Second, syntactic structures exhibit branching bias. Previous studies have not quantitatively examined how this bias affects the relationship between structural and sequential distances, which is crucial for addressing question (ii). Finally, it is well known that PCFGs cannot adequately model natural language syntax without being augmented to relax the context-free independence. However, to answer question (iii) clearly, we need to systematically quantify the differences in the statistical properties between natural language syntax and PCFGs.
In this study, we perform statistical analyses on parse trees to address these three questions, one by one.
@@ 3.1 Datasets and Preprocessing
We primarily use the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000). Its large scale enables highly precise estimation of the statistical quantities.
To verify the robustness of our results, we also conduct the same analyses on two additional corpora. The first is WikiText (Merity et al., 2017), an English corpus from a different domain than BLLIP. The second is the NINJAL Parsed Corpus of Modern Japanese (NPCMJ; National Institute for Japanese Language and Linguistics, 2016), a treebank of Japanese, which is typologically distinct from English.
BLLIP
@@
BLLIP (Charniak et al., 2000) is a large-scale English treebank that provides Penn Treebank (Marcus et al., 1993)-style syntactic annotations for three years’ worth of Wall Street Journal articles. This treebank contains approximately 30 million words and one million sentences in total. We remove the phonological null elements, such as traces of movement and ellipsis, to ensure alignment with WikiText-103 (Merity et al., 2017), which does not include traces.
The number of child nodes differs across nodes in the dataset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees using the Tree.chomsky_normal_form method implemented in the NLTK library (Bird and Loper, 2004). We then apply the Tree.collapse_unary method from the same library to remove the unary branches produced by the former method.
@@
In addition, because a larger number of possible categories requires more data for reliable statistical estimation, we replace part-of-speech (POS) and phrasal tags with a smaller set of categories. Specifically, we refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. The eleven categories used in this study are listed in Appendix A. We apply the same categories to both POS and phrasal tags, as the argument by Lin and Tegmark (2017) treats leaves and internal nodes in the same manner. An example of a preprocessed parse tree is shown in Figure 2.
The statistical quantities analyzed in the present study are defined for the distribution of trees and estimated from the empirical dataset. Because parse trees depend on the chosen annotation scheme, these quantities also depend on the scheme. Nevertheless, we expect that the main results, such as whether the decay is exponential or power-law, will remain unchanged as long as the annotation scheme is reasonable.
To verify this, we also perform the same analyses under two alternative settings: trees without binarization and the phrasal-tag-only setting, in which we distinguish phrasal tags from POS tags and focus only on the former. Appendix B shows that the results for both settings are qualitatively similar to those presented in the main text, that is, the answers to questions (i), (ii), and (iii) are negative.
WikiText
To examine whether our conclusions hold beyond BLLIP, we perform the same analyses on WikiText-103 (wikitext-103-raw-v1 split) (Merity et al., 2017), which consists of English Wikipedia articles that preserves the original Wikipedia markup and contains over 100 million tokens and approximately 3.9 million sentences. We implement a rule-based detokenizer and segment sentences with a spaCy sentencizer (Honnibal et al., 2020). Then, we extract sentences up to 40 words (resulting in approximately 3.4 million sentences) and apply the Berkeley Neural Parser (benepar_en3 model) (Kitaev and Klein, 2018; Kitaev et al., 2019) to them to obtain Penn Treebank-style constituency trees. Finally, we binarize the trees and reduce the number of categories using the same method applied to BLLIP.
NPCMJ
We also analyze the Japanese manually-annotated treebank, NPCMJ (National Institute for Japanese Language and Linguistics, 2016). It contains more than 1.4 million tokens and approximately 90,000 sentences. We first omit all of the phonologically null tokens (null pronouns, traces, etc.) and extract sentences of up to 40 words, resulting in approximately 82,000 sentences. Then we binarize all parse trees and systematically rewrite the POS and phrasal tags using a rule-based mapping scheme listed in Appendix A. Since Japanese is a head-final langauge, all trees are binarized in a left-branching direction, in contrast to the right-branching binarization applied to English.